# Estimating Causal Effects for Binary Outcomes Using Per-Decision Inverse Probability Weighting

### v2

@@ Appendix E Drivers of Efficiency Gain
To understand the drivers of the efficiency gain, here we analytically derive the relative efficiency between the proposed method (we call it “pd-EMEE” where “pd” stands for per-decision) and the EMEE method [11] under a set of simplifying assumptions. In particular, we assume that individuals are always available ($I_{it}\equiv 1$), the randomization probability $p_{t}(H_{it})$ is a constant $p$, the moderator $S_{it}=1$ (i.e., interested in the fully marginal effect), and $\tilde{p}_{t}(S_{it})$ is chosen to be equal to $p$. Furthermore, we consider the case $T=1$ with $\Delta\geq 1$ so for each individual there is only one proximal outcome $Y$ but $\Delta$ treatment occasions and $\Delta$ sub-outcomes $R_{2},\ldots,R_{\Delta+1}$. We assume that all the sub-outcomes are independent with all the treatments, the $\Delta$ sub-outcomes for the same individual are independent with each other, and the sub-outcomes are identically distributed: $R_{t}$ follows Bernoulli($q$) distribution for some $q$.
@@
We derive in Section E.1 that under these simplifying assumptions, the asymptotic relative efficiency between the two estimators is (AVar denoting asymptotic variance)
@@
The relative efficiency (44) equals 1 if $\Delta=1$ and increases with $\Delta$. When $\Delta>1$, the relative efficiency increases with $p$ and $q$ (Figure 4). This confirms our intuition: When $p$ is larger, additional treatments are more likely to occur in the proximal outcome window; when $q$ is larger, sub-outcomes are more likely to be equal to 1 in the proximal outcome window; when $\Delta$ is larger, additional treatments are more likely to occur in the proximal outcome window and sub-outcomes are more likely to be equal to 1 in the proximal outcome window. In all these situations, pd-EMEE can achieve a greater reduction on the terms in the IP weight product, thus improving efficiency.
Figure 4: Relative efficiency (44) against $p$, $q$, and $\Delta$. In each of the first two plots, the three lines for $\Delta=1$ are all flat and overlap with each other.
E.1 Proof of Equation (44)
Suppose all the simplifying assumptions stated in Section E hold.
@@ E.1 Proof of Equation (44)
$=(1-(1-q)^{\Delta})^{-2}\cdot[E\{-e^{-A_{1}\beta}A_{1}(A_{1}-p)\}]^{-2}\cdot E\{e^{-2A_{1}\beta}(A_{1}-p)^{2}\}\cdot L,$ (45)
@@
$=(1-(1-q)^{\Delta})^{-1}\cdot[E\{-e^{-A_{1}\beta}A_{1}(A_{1}-p)\}]^{-2}\cdot E\{e^{-2A_{1}\beta}(A_{1}-p)^{2}\}/(1-p)^{\Delta-1}.$ (46)
Finally, combining (45) and (46) yields (44). This completes the proof.
Appendix F Proof of Equation (10) in Simulation

### v3

@@ E.2 A Theoretical Analysis of the Relative Efficiency
To understand the drivers of the efficiency gain, here we analytically derive the relative efficiency between the proposed method (we call it “pd-EMEE” where “pd” stands for per-decision) and the EMEE method (Qian and others, 2021) under a set of simplifying assumptions. In particular, we assume that individuals are always available ($I_{it}\equiv 1$), the randomization probability $p_{t}(H_{it})$ is a constant $p$, the moderator $S_{it}=1$ (i.e., interested in the fully marginal effect), and $\tilde{p}_{t}(S_{it})$ is chosen to be equal to $p$. Furthermore, we consider the case $T=1$ with $\Delta\geq 1$ so for each individual there is only one proximal outcome $Y$ but $\Delta$ treatment occasions and $\Delta$ sub-outcomes $R_{2},\ldots,R_{\Delta+1}$. We assume that all the sub-outcomes are independent with all the treatments, the $\Delta$ sub-outcomes for the same individual are independent with each other, and the sub-outcomes are identically distributed: $R_{t}$ follows Bernoulli($q$) distribution for some $q$.
@@
We derive in Section E.2.1 that under these simplifying assumptions, the asymptotic relative efficiency between the two estimators is (AVar denoting asymptotic variance)
@@
The relative efficiency (E.1) equals 1 if $\Delta=1$ and increases with $\Delta$. When $\Delta>1$, the relative efficiency increases with $p$ and $q$ (Figure 6). This confirms our intuition: When $p$ is larger, additional treatments are more likely to occur in the proximal outcome window; when $q$ is larger, sub-outcomes are more likely to be equal to 1 in the proximal outcome window; when $\Delta$ is larger, additional treatments are more likely to occur in the proximal outcome window and sub-outcomes are more likely to be equal to 1 in the proximal outcome window. In all these situations, pd-EMEE can achieve a greater reduction on the terms in the IP weight product, thus improving efficiency.
Figure 6: Relative efficiency (E.1) against $p$, $q$, and $\Delta$. In each of the first two plots, the three lines for $\Delta=1$ are all flat and overlap with each other.
E.2.1 Proof of Equation (E.1)
Suppose all the simplifying assumptions stated in Section E.2 hold.
@@
$=(1-(1-q)^{\Delta})^{-2}\cdot[E\{-e^{-A_{1}\beta}A_{1}(A_{1}-p)\}]^{-2}\cdot E\{e^{-2A_{1}\beta}(A_{1}-p)^{2}\}\cdot L,$ (E.2)
@@
$=(1-(1-q)^{\Delta})^{-1}\cdot[E\{-e^{-A_{1}\beta}A_{1}(A_{1}-p)\}]^{-2}\cdot E\{e^{-2A_{1}\beta}(A_{1}-p)^{2}\}/(1-p)^{\Delta-1}.$ (E.3)
Finally, combining (E.2) and (E.3) yields (E.1). This completes the proof.
Appendix F Connection to the Literature
F.1 Connection between Per-Decision IPW (Our Approach) and the Per-Decision Importance Sampling
We describe a related IPW technique in the literature, the per-decision importance sampling (PDIS) in offline reinforcement learning Precup (2000), and discuss how our per-decision IPW approach connects to and differs from PDIS. To make connections to our approach, we describe a simple example of PDIS using a setting similar to MRT. See Liu and others (2020, Section 2) for the general form of PDIS. Suppose there are $\Delta$ decision points, and at each decision point a treatment is randomized with randomization probability $p_{t}$ and a short-term outcome $R_{t+1}$ is observed. Let a distal outcome (an outcome of interest) be $Y:=\sum_{t=1}^{\Delta}R_{t+1}$. Suppose the goal is to estimate $E\{Y(a,\bar{0}_{\Delta-1})\}$. Then we have
$E\{Y(a,\bar{0}_{\Delta-1})\}$ $=E\left[\left\{\frac{\mathbbm{1}(A_{1}=a)}{ap_{1}+(1-a)(1-p_{1})}\prod_{s=2}^{\Delta}\frac{\mathbbm{1}(A_{s}=0)}{1-p_{s}}\right\}\sum_{t=1}^{\Delta}R_{t+1}\right]$ (F.1)
$=E\left[\sum_{t=1}^{\Delta}\left\{\frac{\mathbbm{1}(A_{1}=a)}{ap_{1}+(1-a)(1-p_{1})}\prod_{s=2}^{t}\frac{\mathbbm{1}(A_{s}=0)}{1-p_{s}}\right\}R_{t+1}\right],$ (F.2)
where (F.1) is the identification result using the usual IPW and (F.2) is the identification result using PDIS. The intuition behind (F.2) is that earlier outcomes cannot depend on later treatments. The same intuition is used in our approach to derive the identification result for proximal outcomes satisfying the maximum property; i.e.,
$=E\left(E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{1-p_{j}\left(H_{j}\right)}\right]^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{t+s}=0)}Y_{t,\Delta}\mid A_{t}=a,H_{t},I_{t}=1\right\}\mid S_{t},I_{t}=1\right).$ (F.3)
The difference between our approach and PDIS is that (i) the PDIS literature typically considers the outcome to be a sum of sub-outcomes whereas we considers a binary outcome that is the maximum of sub-outcomes, (ii) the PDIS literature considers a single outcome at the end whereas we considered the setting where the main outcome itself is longitudinal (i.e., we consider $Y_{t,\Delta}$ for $1\leq t\leq T$), and (iii) the PDIS identification result (F.2) is solely based on IPW, whereas our identification result (F.3) is a combination of IPW and standardization.
F.2 Connection between Our Estimating Equation and the Literature
Our estimating function is defined as
$U_{i}(\alpha,\beta)=\sum_{t=1}^{T}I_{it}e^{-A_{it}S_{it}^{T}\beta}\epsilon_{it}(\alpha,\beta)M_{it}W_{it}\begin{bmatrix}g(H_{it})\\ \{A_{it}-\tilde{p}_{t}(S_{it})\}S_{it}\end{bmatrix}.$ (F.4)
Besides the per-decision IP weight, $W_{it}$, which was explained in the previous subsection, our estimating function $U_{i}(\alpha,\beta)$ resembles several existing estimating functions in the literature: the multiplicative structural nested mean model (MSNMM) (Robins, 1994), the estimator for causal excursion effect (EMEE) (Qian and others, 2021), and the weighted and centered least squares (WCLS) estimator (Boruvka and others, 2018).
with $\tilde{\epsilon}_{it}(\alpha,\beta)=Y_{it,1}-g(H_{it}^{T}\alpha-\{A_{it}-\tilde{p}_{t}(S_{it})\}S_{it}^{T}\beta$. Our estimating function (F.4) adopts the weighting (by $M_{it}$) and the centering of the treatment (by $\tilde{p}_{t}(S_{it})$), which were used to (i) estimate a marginal causal effect instead of a fully conditional causal effect, and (ii) make the estimator robust to misspecified control variables (the $g_{t}(H_{it})$ part).
To extend the WCLS estimating function to binary outcome, the EMEE estimator (Qian and others, 2021) used a blipping-down technique of MSNMM developed by Robins (1994), i.e., pre-multiplying the residual term $\epsilon_{it}(\alpha,\beta)$ by a negative causal effect (to remove the causal effect, i.e., “blip” the outcome down to its counterpart under no treatment). We borrowed this idea from EMEE. This is where the term $e^{-A_{it}S_{it}^{T}\beta}$ comes from. Note that the original blipping-down technique Robins (1994) depends on an exponent representing the fully conditional causal effect (such as $e^{-A_{it}f(H_{it})^{T}\theta}$), but we used a blipping-down factor that represents the marginal causal effect because this is the causal effect of interest in our case.
The maximum property may resemble the structure of a discrete-time survival outcome, and may raise the question of whether the proximal outcome should be treated as a survival outcome instead of a binary one. We believe the choice of the model and outcome type should depend on the scientific question. For some interventions, the goal is so that a desirable event, such as self-monitoring in Drink Less and activity bout in HeartStep, occurs within a reasonable time window but not necessarily as soon as possible. For such settings, the causal effect on the binary outcome will contain relevant information for assessing and improving the intervention. For other settings, the goal might be to reduce the risk of a dangerous or harmful behavior as soon as possible, in which case the causal effect on the survival outcome might be more relevant. An example of the latter is problem anger prevention (Metcalf and others, 2022), where the intervention is to reduce the risk of aggressive behavior once anger is detected.
Appendix G Technical Details for the pd-EMEE2 Estimator
In Section 4 we proposed an estimator $\tilde{\beta}$ that leverages the projection idea in semiparametric efficiency theory to further improve the efficiency. We present the technical details for the estimator derivation and the asymptotic theory here.
G.1 Derivation of $\tilde{\beta}$
Here we derive the analytic form of $\sum_{u=1}^{T}[E\{\xi_{i}(\beta)\mid H_{iu},A_{iu}\}-E\{\xi_{i}(\beta)\mid H_{iu}\}]$ mentioned in Section 4, which directly leads to the estimating equation for $\tilde{\beta}$.
The following working assumption facilitates a closed form projection term.
Assumption 8 (Absence of delayed effect)
For any $(t,u)$ pair such that $1\leq u<t\leq T$, $\xi_{it}(\beta)$ is independent of $A_{iu}$ given $H_{iu}$.
The following theorem gives the form of $\sum_{u=1}^{T}[E\{\xi_{i}(\beta)\mid H_{iu},A_{iu}\}-E\{\xi_{i}(\beta)\mid H_{iu}\}]$.
Theorem 8
Under Assumption 8 and Assumptions 14, we have
$~{}~{}~{}~{}\xi_{i}(\beta)-\sum_{u=1}^{T}[E\{\xi_{i}(\beta)\mid H_{iu},A_{iu}\}-E\{\xi_{i}(\beta)\mid H_{iu}\}]$
$=\sum_{t=1}^{T}I_{it}e^{-A_{it}S_{it}^{T}\beta}M_{it}\{A_{it}-\tilde{p}_{t}(S_{it})\}S_{it}\bigg{[}Y_{it,\Delta}W_{it}-E(Y_{it,\Delta}W_{it}|H_{it},A_{it})$
$~{}~{}~{}~{}-\sum_{u=t+1}^{t+\Delta-1}\left\{E(Y_{it,\Delta}W_{it}|H_{iu},A_{iu})-E(Y_{it,\Delta}W_{it}|H_{iu})\right\}\bigg{]}$
$+\sum_{t=1}^{T}I_{it}\tilde{p}_{t}(S_{it})\{1-\tilde{p}_{t}(S_{it})\}S_{it}\left\{e^{-S_{it}^{T}\beta}E(Y_{it,\Delta}W_{it}|H_{it},A_{it}=1)-E(Y_{it,\Delta}W_{it}|H_{it},A_{it}=0)\right\}.$ (G.1)
G.1.1 Proof of Theorem 8
We have
$~{}~{}~{}~{}\xi_{i}(\beta)-\sum_{u=1}^{T}[E\{\xi_{i}(\beta)\mid H_{iu},A_{iu}\}-E\{\xi_{i}(\beta)\mid H_{iu}\}]$
$=\sum_{t=1}^{T}\left[\xi_{it}(\beta)-\sum_{u=t}^{t+\Delta-1}\left\{E(\xi_{it}(\beta)|H_{iu},A_{iu})-E\{\xi_{it}(\beta)|H_{iu}\}\right\}\right],$ (G.2)
because when $u<t$, $E(Y_{it,\Delta}W_{it}|H_{iu},A_{iu})-E(Y_{it,\Delta}W_{it}|H_{iu})=0$ from Assumption 8; when $u\geq t+\Delta$, $\xi_{t}(\beta)$ is deterministic from $H_{u}$.
When $u=t$, $E\{\xi_{it}(\beta)|H_{iu}\}$ is calculated as following:
$E\{\xi_{it}(\beta)|H_{iu}\}$ $=I_{it}S_{it}E\left(e^{-A_{it}S_{it}^{T}\beta}M_{it}\{A_{it}-\tilde{p}_{t}(S_{it})\}Y_{it,\Delta}W_{it}|H_{it}\right)$
$=I_{it}e^{-S_{it}^{T}\beta}\frac{\tilde{p}_{t}(S_{it})}{p_{t}(H_{it})}\{1-\tilde{p}_{t}(S_{it})\}S_{it}E(Y_{it,\Delta}W_{it}|H_{it},A_{it}=1)p_{t}(H_{it})$
$-I_{it}\frac{1-\tilde{p}_{t}(S_{it})}{1-p_{t}(H_{it})}\tilde{p}_{t}(S_{it})S_{it}E(Y_{it,\Delta}W_{it}|H_{it},A_{it}=0)(1-p_{t}(H_{it}))$ (G.3)
$=I_{it}S_{it}\tilde{p}_{t}(S_{it})\{1-\tilde{p}_{t}(S_{it})\}\left\{e^{-S_{it}^{T}\beta}E(Y_{it,\Delta}W_{it}|H_{it},A_{it}=1)-E(Y_{it,\Delta}W_{it}|H_{it},A_{it}=0)\right\}.$ (G.4)
(G.3) follows from the fact that $e^{-A_{it}S_{it}^{T}\beta}$, $M_{it}$ and $A_{it}-\tilde{p}_{t}(S_{it})$ are $(H_{it},A_{it})$-measurable. (G.4) follows from rearranging items in (G.3). Therefore, we also have
$E(\xi_{it}(\beta)|H_{iu},A_{iu})=I_{it}e^{-A_{it}S_{it}^{T}\beta}M_{it}\{A_{it}-\tilde{p}_{t}(S_{it})\}S_{it}E(Y_{it,\Delta}W_{it}|H_{it},A_{it})$ (G.5)
Similarly, when $t<u\leq t+\Delta-1$, $Y_{it,\Delta}$ and $W_{it}$ are not $(H_{u},A_{u})$-measurable, and $E(Y_{it,\Delta}W_{it}|H_{iu},A_{iu})-E(Y_{it,\Delta}W_{it}|H_{iu})$ can be written as
$E(Y_{it,\Delta}W_{it}|H_{iu},A_{iu})$ $=I_{it}e^{-A_{it}S_{it}^{T}\beta}M_{it}\{A_{it}-\tilde{p}_{t}(S_{it})\}S_{it}E(Y_{it,\Delta}W_{it}|H_{iu},A_{iu})$ (G.6)
and
$E(Y_{it,\Delta}W_{it}|H_{iu})$ $=I_{it}e^{-A_{it}S_{it}^{T}\beta}M_{it}\{A_{it}-\tilde{p}_{t}(S_{it})\}S_{it}E(Y_{it,\Delta}W_{it}|H_{iu})$ (G.7)
By plugging (G.4), (G.5), (G.6) and (G.7) in (G.2) we have
$~{}~{}~{}~{}\xi_{i}(\beta)-\sum_{u=1}^{T}[E\{\xi_{i}(\beta)\mid H_{iu},A_{iu}\}-E\{\xi_{i}(\beta)\mid H_{iu}\}]$
$=$ $\sum_{t=1}^{T}I_{it}e^{-A_{it}S_{it}^{T}\beta}M_{it}\{A_{it}-\tilde{p}_{t}(S_{it})\}S_{it}\left[Y_{it,\Delta}W_{it}-\sum_{u=t+1}^{t+\Delta-1}\left\{E(Y_{it,\Delta}W_{it}|H_{iu},A_{iu})-E(Y_{it,\Delta}W_{it}|H_{iu})\right\}\right]$
$-\sum_{t=1}^{T}I_{it}e^{-A_{it}S_{it}^{T}\beta}M_{it}\{A_{it}-\tilde{p}_{t}(S_{it})\}S_{it}E(Y_{it,\Delta}W_{it}|H_{it},A_{it})$
$+\sum_{t=1}^{T}I_{it}S_{it}\tilde{p}_{t}(S_{it})\{1-\tilde{p}_{t}(S_{it})\}\{e^{-S_{it}^{T}\beta}E(Y_{it,\Delta}W_{it}|H_{it},A_{it}=1)-E(Y_{it,\Delta}W_{it}|H_{it},A_{it}=0)\}.$
Rearranging terms and we get (G.1).
G.2 Proof of Theorem 2
Due to Theorem 5.1 in Cheng and others (2023), it suffices to show that $\tilde{U}_{i}(\beta,\mu)$ is globally robust with respect to the nuisance parameter $\mu$; i.e., $E\{\tilde{U}_{i}(\beta^{*},\mu)\}=0$ for all $\mu$.
We have
$\tilde{U}_{i}(\beta,\mu)=I_{1}(\beta)+I_{2}(\beta,\mu)+I_{3}(\beta,\mu),$
where
$I_{1}(\beta)$ $=\sum_{t=1}^{T}I_{it}e^{-A_{it}S_{it}^{T}\beta}M_{it}\{A_{it}-\tilde{p}_{t}(S_{it})\}S_{it}Y_{it,\Delta}W_{it},$
$I_{2}(\beta,\mu)$ $=-\sum_{t=1}^{T}I_{it}e^{-A_{it}S_{it}^{T}\beta}M_{it}\{A_{it}-\tilde{p}_{t}(S_{it})\}S_{it}$
×$~{}~{}~{}~{}\times\sum_{u=t+1}^{t+\Delta-1}\Big{\{}\mu_{u-t}(H_{iu},A_{iu})-p_{u}(H_{iu})\mu_{u-t}(H_{iu},1)-(1-p_{u}(H_{iu}))\mu_{u-t}(H_{iu},0)\Big{\}},$
$I_{3}(\beta,\mu)$ $=\sum_{t=1}^{T}I_{it}S_{it}\bigg{[}-e^{-A_{it}S_{it}^{T}\beta}M_{it}\{A_{it}-\tilde{p}_{it}(S_{it})\}\mu_{0}(H_{it},A_{it})+\tilde{p}_{t}(S_{it})\{1-\tilde{p}_{t}(S_{it})\}\Big{\{}e^{-S_{it}^{T}\beta}\mu_{0}(H_{it},1)-\mu_{0}(H_{it},0)\Big{\}}\bigg{]}.$
That $E\{I_{1}(\beta^{*})\}=0$ follows from Lemma 7. That $E\{I_{2}(\beta^{*},\mu)\}=0$ for all $\mu$ follows from the law of iterated expectation and the fact that
$E\{\mu_{u-t}(H_{iu},A_{iu})\mid H_{iu}\}=p_{u}(H_{iu})\mu_{u-t}(H_{iu},1)-(1-p_{u}(H_{iu}))\mu_{u-t}(H_{iu},0)\quad\text{for any }\mu.$
That $E\{I_{3}(\beta^{*},\mu)\}=0$ for all $\mu$ follows from the law of iterated expectation and the fact that
$~{}~{}~{}~{}E[e^{-A_{it}S_{it}^{T}\beta}M_{it}\{A_{it}-\tilde{p}_{it}(S_{it})\}\mu_{0}(H_{it},A_{it})\mid H_{it}]$
$=E[e^{-A_{it}S_{it}^{T}\beta}M_{it}\{A_{it}-\tilde{p}_{it}(S_{it})\}\mu_{0}(H_{it},A_{it})\mid H_{it},A_{it}=1]p_{t}(H_{it})$
$~{}~{}~{}+E[e^{-A_{it}S_{it}^{T}\beta}M_{it}\{A_{it}-\tilde{p}_{it}(S_{it})\}\mu_{0}(H_{it},A_{it})\mid H_{it},A_{it}=0]\{1-p_{t}(H_{it})\}$
$=\tilde{p}_{t}(S_{it})\{1-\tilde{p}_{t}(S_{it})\}\Big{\{}e^{-S_{it}^{T}\beta}\mu_{0}(H_{it},1)-\mu_{0}(H_{it},0)\Big{\}}.$
This completes the proof.
Appendix H Generative model
H.1 Generative Model used in Simulation
We set $I_{t}\equiv 1$ and $T=100$. The generative model depends on two parameters, $\Delta$ (length of the proximal outcome window) and $p_{a}$ (constant randomization probability). $A_{t}\in\{0,1\}$ was generated from $\text{Bernoulli}(p_{a})$. A scalar time-varying covariate, $Z_{t}\in\{0,1,2\}$, was generated independently of all previous variables (i.e., $Z_{t}\perp\{Z_{s},A_{s},R_{s+1}:1\leq s\leq t-1\}$) with
$P(Z_{t}=0)=0.5^{-1/2\Delta}/C,\quad P(Z_{t}=1)=1/C,\quad P(Z_{t}=2)=0.5^{1/2\Delta}/C,$ (H.1)
$P(R_{t+1}=0\mid A_{t}=0,H_{t})$ $=0.5^{(1.5-0.5Z_{t})/\Delta},$ (H.2)
$P(R_{t+1}=0\mid A_{t}=1,H_{t})$ $=\frac{1-\{1-0.5^{(1.5-0.5Z_{t})/\Delta}\cdot(3/C\cdot 0.5^{1/\Delta})^{\Delta-1}\}\cdot e^{0.1+0.2Z_{t}}}{(3/C\cdot 0.5^{1/\Delta})^{\Delta-1}}.$ (H.3)
We will show in the next subsection that for $1\leq t\leq T-\Delta$,
$E\{Y_{t,\Delta}(\bar{A}_{t-1},a_{t},\bar{0}_{\Delta-1})\mid H_{t}(\bar{A}_{t-1}),I_{t}(\bar{A}_{t-1})=1\}=\{1-0.5^{(\Delta+0.5-0.5Z_{t})/\Delta}\cdot(3/C)^{\Delta-1}\}\cdot e^{a_{t}(0.1+0.2Z_{t})}.$ (H.4)
There is substantial variability in $R_{t+1}$ and $Y_{t,\Delta}$. When we vary $\Delta$ from 2 to 10 and vary $Z_{t}\in\{0,1,2\}$, (H.2) varies from 0.59 to 0.97, (H.3) varies from 0.38 to 0.80, and (H.4) varies from 0.41 to 0.58 when setting $a_{t}=0$ and from 0.57 to 0.80 when setting $a_{t}=1$.
We constructed this generative model based on three considerations. First, in order to fairly assess how the efficiency improvement is affected by $\Delta$ and $p_{a}$, we needed $\text{CEE}(Z_{t})$ to not depend on $\Delta$ or $p_{a}$. Thus, the distribution of the sub-outcome $R_{t}$ would depend on $\Delta$ because there are $\Delta$ sub-outcomes in $Y_{t,\Delta}$. Second, we wanted to construct a complicated $E\{Y_{t,\Delta}(\bar{A}_{t-1},0,\bar{0}_{\Delta-1})\mid H_{t},I_{t}=1\}$ to demonstrate the robustness of the estimator against misspecified working models. Third, we wanted $Z_{t}$ to depend on $\Delta$ as little as possible. In a real application, $Z_{t}$ (a covariate) would not depend on $\Delta$ (a choice in the analysis). However, we found it difficult to achieve while keeping the form of $\text{CEE}(Z_{t})$ simple. The current generating distribution of $Z_{t}$ (H.1) is the least dependent on $\Delta$ that we came up with, and the three probabilities in (H.1) are all very close to 1/3 for $\Delta\geq 2$.
H.2 Proof of Equation H.4